OpenStudy (lena772):

Is this right? Find d^2y/dx^2 in terms of x and y for y^2 = 4x: y^2 = 4x 2yy' = 4 yy' = 2 y(y'') + y'(y') = 0 sqrt(4x)y'' + (2/y)^2 = 0 y'' = -(2/y)^2 / sqrt(4x)

OpenStudy (luigi0210):

Oh, I love implicit differentiation, but I've never dealt with the second derivative of it ._.

OpenStudy (luigi0210):

But wouldn't it start like this?: \[\LARGE y^2=4x\] \[\LARGE 2yy'=4\] \[\LARGE y'=\frac{2}{y}\] Sorry, I'm rusty on this stuff ._.

OpenStudy (shamil98):

2(y^-1) \[\large y^{-1} + 2y^{-2}y'\] i think

OpenStudy (shamil98):

\[\large y'(2y^{-2}) = y^{-1}\]

OpenStudy (shamil98):

-y^-1*

OpenStudy (shamil98):

Then the simplification.. if im doin it rite

OpenStudy (luigi0210):

This is messy but: \[\LARGE y''= \frac{y(-2)'-y'(2)}{y^2}\] Simplify: \[\LARGE y''=\frac{0-2y'}{y^2}\] \[\LARGE y''=-\frac{2y'}{y^2}\] Now we plug what we got from y' back in: \[\Huge y''=-\frac{2({\frac{2}{y}})}{y^2}\] And we simplify from here ._.

OpenStudy (shamil98):

\[\large y''(\frac{ 1 }{2y^2 }) = \frac{ -1 }{ y } \] \[\large y'' = \frac{ -1 }{ y } \times 2y^2\] \[\large y'' = -2y\] i think i did something wrong O_O

OpenStudy (luigi0210):

We learn through mistakes :3